# VecTrans: LLM Transformation Framework for Better Auto-vectorization on High-performance CPU

### v1

@@ 3.1. Source 1: Self-Feedback
Figure 3. The compiler provides feedback through lexical/syntactic errors, semantic unit test results, and vectorization insights via Rpass flags. This multi-source diagnostic evaluates code validity and vectorization potential.
3.2. Source 2: Compiler Feedback
@@
Compiler feedback information consists of three parts as shown in Figure 3. In the first part, it conducts a compilation-error check on the optimized code, capitalizing on the error-related information provided by the compiler. In the second part, it performs a semantic-equivalence analysis by comparing the source code and the optimized code through unit test. In the third part, the Rpass and Rpass-analysis parameters of the compiler output vectorized information about the optimized code, including the positions of the loops that were successfully vectorized, the positions of the loops that were not successfully vectorized, and the reasons why the vectorization was not successful. These three parts of information can help the LLM determine whether the current code is optimized and fully vectorized while ensuring semantic correctness.
Figure 4. Verification feedback ensures code correctness through dual mechanisms. Unit testing verifies semantic equivalence by comparing execution outputs, though constrained by input coverage limitations. Formal verification complements this approach by mathematically proving behavioral equivalence through specification-based analysis, establishing multi-layered correctness guarantees.
3.3. Source 3: Verification Feedback
@@
The verification feedback is divided into unit test information and formal verification information shown in Figure 4:
1. Unit test: This tests the syntax and semantic correctness of the code. The generated optimized code must first pass compilation and ensure that its semantics are consistent with the source code. In this phase, the LLM is used to generate unit test code, which includes generating simple inputs and validating outputs. Unit tests allow the correctness of the optimized code to be verified within the scope of the covered inputs. To ensure efficiency, unit test code is generated only once at the beginning and reused in subsequent iterations. To ensure the correctness of the unit test code, it is validated by replacing the optimized code. First, the original function replaces the optimized function, and the output of the optimized function must match the original function, as the two functions are identical at this point. If the return value is ’void’, an empty function is used to replace the optimized code. If the return value is not ’void’, the return values of the original code are bitwise negated. At this point, the outputs of the original and optimized functions must differ; otherwise, the unit test code is considered incorrect. There is a flaw in this approach, as if the original function is empty, a false positive might occur, but this situation is rare, so it is ignored here.
2. Formal verification: After the optimized code passes the unit tests, it must undergo formal verification. Alive2(10.1145/3453483.3454030, ) is used as the formal verification tool. Both the original and optimized functions are compiled into IR and then verified using Alive2. If the formal verification is successful, the code moves to performance testing. If it fails, the process loops back with the addition of error information.
3.4. Combination of Feedbacks
@@
In the refine phase, a comprehensive set of inputs, namely the source code, the current state of the optimized code, the feedback information, and relevant prior knowledge, is integrated. The prior knowledge encompasses concepts such as loop splitting, loop reordering, instruction reordering, conditions for using temporary variables, along with simple illustrative examples like Instruction Splitting (which requires independence between instructions) and Iteration Range Splitting (where the loop should contain phase-dependent patterns).
This phase capitalizes on the feedback from the previous feedback phase. The LLM receives as inputs the feedback data, the source code, and the optimized code. Leveraging this information and the prior knowledge, the model makes informed decisions to update the optimized code. The output of this process is the refined optimized code, which is carefully crafted to maximize performance and adhere to programming standards.
A key aspect of this phase is the constraint of generating only the optimized code, which is encapsulated with specific identifiers. This not only streamlines the output but also enables easier post-processing and integration. By decoupling the feedback and iteration phases, we enhance the LLM’s performance. It allows for more focused processing within each phase, facilitating better organization of the overall optimization process. This, in turn, makes the entire procedure more conducive to automation, as the clear separation of tasks and well-defined input-output relationships simplify the implementation of automated workflows.
3.5. Optimized Case
The framework has the ability to handle complex functions from real-world applications(Skia), as shown in Figure 5. The temporary array is used to temporarily store intermediate results, and loop splitting is used to separate the parts with loop dependencies, so that the remaining loops without dependencies can be vectorized successfully.
@@ 3.5. Optimized Case
(a) Original code

### v2

@@ III-A Self-Feedback
Figure 3: The compiler provides feedback through lexical/syntactic errors, semantic unit test results, and vectorization insights via Rpass flags.
III-B Compiler Feedback
@@
Compiler feedback information consists of three parts as shown in Figure 3. In the first part, it conducts a compilation error check on the optimized code, capitalizing on the error-related information provided by the compiler. In the second part, it performs a semantic-equivalence analysis by comparing the source code and the optimized code through unit tests. In the third part, the Rpass and Rpass-analysis parameters of the compiler output vectorized information about the optimized code, including the positions of the loops that were successfully vectorized, the positions of the loops that were not successfully vectorized, and the reasons why the vectorization was not successful. These three types of information enable the LLM to assess whether the code is optimized and fully vectorized while preserving semantic correctness.
Figure 4: Verification combines: Unit Test - Output-equivalence testing across representative inputs; Formal verification - Specification-based analysis of behavioral equivalence.
III-C Verification Feedback
@@
The verification feedback is divided into unit test information and formal verification information, shown in Figure 4:
Unit testing: The unit testing phase verifies both syntactic correctness (via compilation) and semantic equivalence between optimized and original code. The framework leverages LLM-generated tests to validate semantic preservation through I/O comparisons. Tests are generated once during initialization and reused across iterations to minimize overhead. Test validity is ensured via two-step verification: (1) Baseline Equivalence: The optimized function is replaced with the original to confirm identical outputs, ensuring test reliability. (2) Test Sensitivity Check: For non-void returns, the original function is modified to return bitwise-negated results, replacing the optimized code to enforce a controlled mismatch. For void returns, the optimized code is replaced with an empty implementation to induce behavioral divergence. This approach reliably detects invalid test suites, with rare false positives only occurring for empty original functions—an edge case excluded due to negligible practical impact.
Formal verification: Following unit testing, the optimized code undergoes formal verification using Alive2[18], where both original and optimized functions are compiled to IR and analyzed. Crucially, this verification occurs prior to vectorization, ensuring minimal divergence between IR versions and thus simplifying the verification process. Alive2 may produce four distinct outcomes: semantic equivalence (proceeding to performance testing), semantic mismatch, timeout, or verification error (triggering corrective feedback). Successful verification advances the pipeline, while failures initiate iterative refinement with diagnostic error information.
III-D Combination of Feedbacks
@@
In the refine phase, a comprehensive set of inputs is integrated, including the source code, the optimized code, feedback information, and relevant prior knowledge. This prior knowledge covers optimization techniques such as loop splitting, loop reordering, instruction reordering, and the use of temporary variables, along with simple illustrative patterns like instruction splitting (which requires independence between instructions) and iteration range splitting (applicable when loops exhibit phase-dependent behaviors).
This phase builds upon the feedback collected in the previous stage. The LLM takes the feedback data, source code, and current optimized code as input, and, guided by the embedded prior knowledge, iteratively refines the optimized code. The result is an improved code version, explicitly designed to enhance performance while preserving correctness and adhering to established coding conventions.
A notable characteristic of this phase is the strict constraint that the LLM generates only the optimized code, encapsulated with specific identifiers. This design choice simplifies downstream post-processing and facilitates seamless integration. Furthermore, by decoupling the feedback and refinement phases, the framework improves LLM efficiency and maintains a clear separation of concerns. This modular design enhances the scalability and automation potential of the optimization workflow by enforcing well-defined input-output interfaces for each phase.
IV Evaluation
IV-A Experimental Methodology
System Configuration. We conduct experiments on Linux-based platforms with the following two specifications:
  • Arm Platform: Kunpeng 920, 128-core@2.6GHz
  • X86 Platform: Intel Xeon Gold 6132, 28-core@2.6GHz
We use GCC 14.2.0, Clang 20.0.0, BiSheng Compiler 3.2.0.1[19], and Intel(R) oneAPI DPC++/C++ Compiler (ICX) 2023.2.0, with compile flag -O3 -ffast-math and vectorization diagnostics enabled via -Rpass=loop-vectorize -Rpass-analysis=loop-vectorize. We utilize the Deepseek-V3 API and conduct comparative analyses against the Qwen2.5-Coder-32B-Instruct and Qwen2.5-72B-Instruct models.
Benchmarks. We evaluate a wide range of functions from TSVC_2[20] and real application function from Skia[21]. The TSVC_2 (Test Suite for Vectorization Compilers) benchmark suite systematically evaluates compiler capabilities in automatic vectorization through 149 test functions representing typical computational patterns. Former results reveal that 50 functions (approximately 33.6%) resist effective vectorization by BiSheng Compiler[19] and mainstream compilers, primarily due to complex data dependencies, irregular memory access patterns, or unstructured control flows. The classification of these functions is shown in Table I. These selected cases serve as critical evaluation targets, exposing limitations in current vectorization techniques while providing a standardized test suite for analyzing compiler optimization boundaries and guiding algorithmic improvements.
Metrics. We assess framework effectiveness through three key metrics: (1) Coverage: Percentage of compiler-unvectorizable functions successfully transformed (primary success metric); (2) Speedup: Performance gain of vectorized functions (benefit metric); (3) Iteration Cost: Optimization rounds required (deployment cost metric).
Parameter Setup. The LLM selected for the experiment is Deepseek-V3[22]. The maximum number of iteration rounds is set to 20. Set the maximum number of tokens to 4096 and retain the default values for other parameters.
TABLE I: The classification of functions that cannot be vectorized in TSVC_2.
Category Function name
unable to identify reduction variable (33) s116, s123, s126, s141, s161, s211, s212, s1213, s221, s222, s231, s232, s233, s2233, s235, s2251, s256, s258, s261, s275, s277, s291, s292, s2111, s318, s3110, s3112, s321, s322, s323, s341, s342, s343
unable to identify array bounds (2) s123, s1161
unable to determine number of loop iterations (5) s161, s277, s332, s481, s482
instructions (e.g., call instruction) cannot be vectorized (2) s31111, s451
loop contains switch statements (1) s442
Figure 5: VecTrans delivers a geometric mean speedup of 1.77×, marginally outperforming LLM-Vectorizer’s 1.74×. Shaded bars denote vectorization failures. The evaluation covers 29 functions, representing the union of vectorizable cases. Both frameworks use Deepseek-V3, with LLM-Vectorizer re-evaluated on Neon and Deepseek-V3 for consistency.
IV-B Speedup
Our evaluation demonstrates that VecTrans, when integrated with the BiSheng Compiler[19], successfully vectorizes 24 functions (23 from TSVC and 1 from Skia) in our benchmark suite. As shown in Figure 5, VecTrans delivers a 1.77x geometric mean speedup over the -O3 baseline, peaking at 5.21x for function s293. When evaluated under ICX, our technique yields a 1.2× geometric mean speedup.
However, results reveal two limitations: (1) two functions exhibit performance degradation (up to 19% slowdown), and (2) four functions show negligible improvement (speedup <5%). A representative example is function s442, which achieves only a 0.81× speedup after transformation. This degradation stems from additional computations introduced during vectorization. Specifically, s442 contains a switch statement, and to enable automatic vectorization, the transformation replaces each branch with a condition-masked computation. For example, the transformed code computes every possible branch result and selects the correct one via condition multipliers (e.g., (condition) $\times$ (statement)). Microarchitectural analysis reveals three primary causes for these suboptimal cases: (1) memory overhead from temporary array allocations, (2) increased cache miss rates due to non-contiguous memory access patterns, and (3) pipeline stalls caused by vector instruction scheduling conflicts. These findings highlight the critical trade-offs between vectorization potential and architectural constraints that must be carefully balanced in automatic vectorization frameworks.
IV-C Coverage
Our framework demonstrates significant improvements in vectorization coverage, successfully vectorizing 23 of 50 TSCV functions (46%), as well as complex real-world cases from Skia. This substantially outperforms LLM-Vectorizer’s 28.8% coverage on Neon ISA targets, highlighting the impact of limited architecture-specific training data in existing approaches.
Figure 5 and Table III present the vectorization success rates across the categories in Table I. Our framework achieves the highest efficacy (10/13 success) on functions previously unvectorizable due to unsafe memory dependencies. Reduction variable identification posed greater challenges, resulting in only 13 out of 33 functions being vectorized. The remaining four categories posed significant optimization barriers—only s1161 was successfully vectorized, with no additional functions from these groups being optimized. Figure 5 also presents speedups for 29 functions, corresponding to the union of those vectorized by VecTrans and LLM-Vectorizer. Notably, LLM-Vectorizer excels in handling conditional branches (right side of Figure 5), particularly in functions s3110 to s342, which compute array maxima or minima with irreducible control flow. These constructs typically resist conventional auto-vectorization and persist even after VecTrans transformations. LLM-Vectorizer addresses this by merging multiple conditional checks into single vector instructions, though parallelism remains constrained, limiting achievable speedups. Conversely, VecTrans demonstrates superior performance on loops with cross-iteration dependencies (e.g., s2233–s281), where it systematically eliminates reducible dependencies through targeted transformations, enabling effective compiler auto-vectorization. While LLM-Vectorizer can occasionally discover dependency-breaking patterns during exploration, its lack of explicit dependency elimination strategies results in less consistent outcomes compared to the principled, dependency-aware approach of VecTrans. Together, the two methods complement each other, collectively broadening the scope of vectorizable code.
Table II shows the methods in which the LLM transformation function enables vectorization. Through these transformations, the compiler can more easily identify whether there is a dependency between loops, so that the compiler can vectorize a function that could not be vectorized originally. Among optimization techniques, loop splitting emerges as the most effective solution for eliminating false dependencies in loop structures. Instruction reordering and temporary variable insertion are employed to enhance compiler dependency analysis, which not only improves false dependency identification but also prevents unsafe memory accesses such as simultaneous array read-write operations. While branch elimination converts conditional jumps into arithmetic expressions, this approach inevitably introduces substantial redundant computations, thus generally offering no performance benefits.
TABLE II: Code Optimization Techniques and Descriptions
Optimization Description
Temporal Variable Introduce short-lived variables to store recurrent computation results, reducing dependency caused by array access.
Branch Elimination Remove conditional statements like switch or if.
Beyond the successfully vectorized cases, the framework fails to transform 27 test functions. Most of these functions lack vectorization potential, and forced vectorization would lead to performance degradation. The failures occur due to:
  • Iteration Limit Exceeded (9 cases). The iteration upper bound is set to 20. If the LLM fails to vectorize the function after 20 iterations, the process terminates, and the original code is preserved.
  • Premature Optimization Assumption (2 cases). In some cases, the LLM disregards compiler-reported vectorization constraints, incorrectly concluding that the code is already optimized. This leads to early termination without meaningful vectorization.
  • Loops with No Vectorization Benefit (15 cases). The LLM identifies loops where forced vectorization offers no performance gain (e.g., due to loop-carried dependencies or conditional branches). Attempting vectorization would introduce redundant memory copies, degrading performance (deoptimization). Thus, the original code is preserved.
  • Semantic Errors (1 case: s481). A single failure (1/51 tests) produced semantically incorrect code despite passing unit tests and formal verification. This stems from insufficient test coverage and limitations in the verification toolchain. Such errors are critical and will be addressed in future work.
More specifically, the framework encounters vectorization challenges primarily due to four key patterns: (1) complex conditional branches (affecting 11 test cases), (2) unresolvable loop-carried dependencies (8 cases), (3) irregular memory access patterns from multi-dimensional arrays (6 cases), and (4) external function calls requiring inlining (2 cases). Among these, conditional branches constitute the dominant limiting factor, while array-induced memory access complexities and loop dependencies present secondary obstacles. The remaining cases involve function calls that necessitate source-level inlining—a transformation currently beyond the framework’s scope.
TABLE III: The coverage and rounds under different configurations. The base model means that the framework is without formal verification, compiler feedback, and unit tests.
Configuration Coverage Geomean speedup Iteration rounds
DeepSeek-V3 / VecTrans $\mathbf{46.2\%}$ 1.77x 8.760
DeepSeek-V3 / LLM-Vectorizer[16] 28.8% 1.74x 13.39
DeepSeek-V3 / base model 17.3% 1.57x 5.412
Qwen2.5-32B / VecTrans 34.6% 1.96x 13.940
Qwen2.5-72B / VecTrans 38.5% 1.87x 12.260
DeepSeek-V3 w/o formal verification 32.7% 1.76x 8.489
DeepSeek-V3 w/o compiler feedback 21.2% 1.74x 5.213
DeepSeek-V3 w/o unit test 25.0% 1.52x 9.656
Figure 6: The graph shows the speedup of functions that can be successfully vectorized when the framework uses different LLMs. Columns with slashes indicate that vectorization cannot be successful. Three LLMs, Qwen2.5-Coder-32B-Instruct[23], Qwen2.5-72B-Instruct[24] and Deepseek-V3, are used here. The speedup is of the final generated optimized function compared to the original function using BishengCompiler at -O3 -ffast-math flag.
IV-D Iteration Rounds
The framework requires 8.76 iterations per function on average, with 9 functions reaching the 20-iteration ceiling. Successful vectorization demonstrates significantly greater optimization efficiency at 4.96 iterations (43.4% reduction from baseline).
Token consumption analysis reveals 2.5k input/600 output tokens per iteration. Given the average 21.9k input/5.3k output tokens per function and DeepSeek-v3’s pricing of \$0.27/M input and \$1.10/M output tokens, the vectorization cost per function averages \$0.012.
IV-E Breakdown Analysis
Table III presents quantitative evidence of component effectiveness through coverage metrics and iteration dynamics across architectural variants. Our ablation study reveals significant performance degradation when disabling individual components: removing compiler feedback causes the most substantial coverage drop (46.2% → 21.2%, -54.1%), followed by unit test exclusion (46.2% → 25.0%, -45.9%), with formal verification removal showing the least impact (46.2% → 32.7%, -29.2%). The baseline configuration (without all three components) achieves only 17.3% coverage, demonstrating their cumulative effectiveness through 2.67× improvement over the base model.
Compiler feedback enhances the framework in three key ways: (1) Helps LLM to determine whether a compilation error occurs. If a compilation error occurs, the optimized function needs to be regenerated. (2) Helps LLM determine whether vectorization is successful. If vectorization has not been performed, LLM needs to analyze whether vectorization opportunities exist and provide conversion suggestions. (3) Assists the LLM in identifying transformations that enable automatic vectorization by the compiler. The compiler gives the loop position that cannot be vectorized and the reason why it cannot be vectorized.
The removal of different components has an impact on the coverage rate and iteration rounds, but has little impact on the acceleration ratio because the code optimization methods used by the same LLM are similar, and the generated code is basically the same. Except that a few functions may reduce the acceleration ratio due to redundant operations, the acceleration ratio remains the same in other cases.
Figure 7: Geometric mean speedup of functions optimized by VecTrans across different configurations. Each bar reports the speedup relative to the O3 baseline of the corresponding compiler and hardware backend.
IV-F Sensitivity Studies
Table III presents the coverage rate and iteration rounds for the three base models. As model parameters decrease, we observe a reduction in coverage (DeepSeek-V3: 46.2%, Qwen2.5-72B: 38.5%, Qwen2.5-32B: 34.6%) and an increase in iteration rounds (8.76, 12.26, and 13.94, respectively). This trend suggests that smaller models exhibit weaker capability, failing to transform some test functions successfully. Consequently, they require more iterations, with some cases failing to produce correct results within the 20-round limit.
Figure 6 shows the speedup of different model generation functions. The speedup of Qwen2.5-Coder-32B-Instruct is 1.96 times, and that of Qwen2.5-72B-Instinct is 1.87 times. The speedup of Deepseek-V3 is 1.77 times. Even for different LLMs, the generated functions are similar, and the speedup is similar. Therefore, this framework may help identify potential compilation optimization points, that is, summarize the characteristics of the generated functions. The optimization direction of automatically vectorized can be found.
Sensitivity of different compilers. Here we explore the impact of different compilers on the optimized functions, and we test the speedup of GCC, Clang, ICX, and BiSheng Compiler[19], respectively. The GCC and Clang versions used here are the latest versions, which represent the best compilers available today. ICX is Intel-optimized for superior CPU/GPU performance on Intel hardware, outperforming GCC/Clang in Intel-centric environments. BiSheng Compiler is a compiler optimized for Kunpeng series chips. GCC, Clang, and BishengCompiler are tested on Kunpeng 920, while ICX is tested on Intel CPU. The Kunpeng 920 uses the Neon instruction set, and Intel CPU employs the AVX instruction set. The speedup of functions compiled using the flag -O3 -ffast-math on different compilers is shown in Figure 7.
The geomean speedup of the optimized function is 1.42 times on Clang, 1.41 times on GCC, and 1.77 times on BiSheng Compiler[19]. The code generated for the s112 function is significantly worse on GCC and Clang than it is on BiSheng Compiler[19] because the latest versions of GCC and Clang can vectorize the original s112 function. In s1161 and s242, the speedup of GCC is lower than that of the other two compilers. In s293, the speedup of Clang is lower than that of the other two compilers. This indicates that different compilers optimize functions differently. However, in most cases, the optimization of these functions is effective for all three compilers, and can achieve good results. The speedup achieved with ICX was 1.2×, which is lower than that of the other three compilers. This discrepancy arises not only from differences in test hardware (Kunpeng 920 utilizing Neon instructions vs. Intel CPUs leveraging AVX) but also because ICX incorporates architecture-specific optimizations for Intel Xeon CPU, resulting in higher vectorization coverage that reduces its geomean speedup gains in this cross-platform comparison.
As shown in Figure 7, the geomean speedup is 2.02 times on the AVX instruction set, 1.77 times on the Neon instruction set, 1.94 times on the SVE instruction set, and on the AVX instruction set, the speedup of the s1161 function can reach 16.8 times. It can be observed that, on different vector instruction sets, there is no obvious regularity in the speedup of the function, but the effect is better on the AVX instruction set and the SVE instruction set. The effect comes from the length of the vectorized register on the one hand, and the vectorized implementation mode on the other hand.
void s481(...) {
for (int nl = 0; nl < iters; nl++){
for (int i = 0; i < LEN_1D; i++){
if (d[i] < (float)0.) return;
a[i] += b[i] * c[i];
(a) Original code that performs updates to a until a negative value is detected in d, exiting immediately afterwards.
void s481_opt(...) {
for (int nl = 0; nl < iters; nl++){
int early_exit = 0;
for (int i = 0; i< LEN_1D; i++)
if (d[i] < (float)0.){
early_exit = 1;
break;
if (early_exit) return;
for (int i = 0; i < LEN_1D; i++)
a[i] += b[i] * c[i];
(b) Incorrect optimization that hoists the negative check, altering the original semantics by potentially skipping valid updates.
Figure 8: Failure case: The transformation incorrectly hoists the termination condition, potentially omitting valid updates that would occur before negative values in the original execution.
IV-G Case Study
IV-G1 Failure Case Analysis
As shown in Figure 8, test case s481 generated erroneous code that passed both unit testing and formal verification. This section analyzes how this test case circumvented correctness verification. First, the test case passed unit testing due to randomly generated inputs. The large language model’s generated test functions predominantly contained positive values in their random inputs, whereas triggering the conditional branch in this function requires negative values in the d-array. Consequently, the unit test inputs failed to cover this scenario, resulting in the function erroneously passing verification. This limitation can be mitigated by constraining the model to generate inputs with improved coverage characteristics. Second, this test case passes Alive2’s verification because Alive2 can only perform bounded loop checking. It operates by unrolling loops in the target function for a fixed number of iterations and verifying the semantics of the unrolled version. Consequently, Alive2’s soundness is constrained by this bounded verification approach, meaning it cannot guarantee correctness under all possible scenarios.
IV-G2 Real World Application
In this section, we use real-world function from Skia[21], as shown in Figure 9, to show how VecTrans works in detial. Skia is an open-source 2D graphics rendering engine developed by Google, extensively utilized in projects such as Chrome, Android, and Flutter. It delivers high-performance support for vector graphics, text rendering, image processing, and GPU-accelerated rendering. In the source code, the presence of loop-carried dependencies in the dst array prevents vectorization. In the optimized code, the temporary array is used to temporarily store intermediate results, and loop splitting is used to separate the parts with loop dependencies, so that the remaining loops without dependencies can be vectorized successfully. After optimization by VecTrans, this function achieves a 1.35x speedup on the Neon instruction set when compiled with BiShengCompiler.
@@ IV-G2 Real World Application
(a) Original implementation with data-dependent control flow